078898:

how to solve.Ruby is visiting a wildlife center to gather information for her term paper. The center has a circular pond with a diameter of 20 meters. What is the approximate area of the pond?

Hero:

Area of a circle is \(A = \pi r^2\) where \(r = \dfrac{d}{2}\) in this case.